# A systematically-selected sample of luminous, long-duration, ambiguous nuclear transients

### v1

@@ 3.1.2 Spectroscopically selected ANTs
ZTF20acvfraq/AT2020adpi/ATLAS20bjzp/Gaia21aid, . First reported by ZTF. The rise time is uncertain due to gaps in the observed light curve, constrained to d d in the rest frame. The initial decline is shallow, before a steep phase and then a second shallow phase. We obtain a Keck/LRIS spectrum from WISeREP. The spectrum is similar to ZTf19aailpwl/AT2019brs: dominated by strong Balmer lines with broad (FWHM km s-1 and narrow components, along with He i and Mg ii with similar profiles. Similarly to ZTF19aailpwl/AT2019brs is the presence of He ii, N iii, and a particularly strong O iii line. These features indicate Bowen fluorescence. Fe ii features can also be seen in the spectrum. A MIR flare is present with a lag of yr at peak. No X-ray observations have been made of this source.
ZTF19aatubsj/AT2019fdr/ATLAS19lkd/PS19dar/Gaia19bsz, . First reported by ZTF. Classified as a TDE in a NLSy1 (Frederick
et al., 2021; Reusch
et al., 2022), and although Pitik
et al. (2022) classify it as a SLSN-IIn, we favour a SMBH-related interpretation. The light curve is not smooth: the rise displays a ‘shoulder’ before peaking after 70 d, while after 70 d of steep decline there is an 80 d plateau before a shallower decline. After 500 days a 0.7 mag rebrightening occurred and the transient has remained close to that level ever since. There is a MIR flare with a lag at peak of d. The spectrum has narrow + broad Balmer emission lines as well as Fe ii emission. There is no clear detection of He ii or Bowen fluorescence lines. A MIR flare is present with a lag of yr at peak. X-ray observations are presented in detail by Reusch
@@
ZTF20aanxcpf/AT2021loi/ATLAS21qje, . First reported by ZTF. This event was analysed by Makrygianni et al. (2023) as a Bowen fluorescence flare. It is the lowest luminosity of our sample ( erg s-1). It has a similar initial light curve shape to the majority of the main sample presented here, but displays a clear rebrightening after d in the rest frame. The spectrum is similar to ZTF19aailpwl/AT2019brs and ZTF20acvfraq/AT2020adpi in that it is dominated by Balmer lines with an unusual, flat-topped shape. There are strong emission lines from He ii, N iii and O iii, indicating Bowen fluorescence and thus a very hot ionizing source. Any Fe ii contribution is much weaker than in some of the Fe-strong events, but here the Balmer continuum is stronger. A MIR flare is present with a lag of yr at peak.
@@
ZTF20abrbeie | AT2021lwx | 198 | 45.9 | 2560a | 250 | - | |||
ZTF19aamrjar | - | 377 | 45.6 | 6200b | 732 | 41.5 | |||
ZTF20abodaps | AT2020afep | 17 | 45.5 | 1540 | 40.7 | ||||
ZTF18aczpgwm | AT2019kn | 125 | 45.2 | 2831c | 245 | 40.6 | |||
ZTF21abxowzx | AT2021yzu | 71 | 45.2 | 1690 | 194 | 40.2 | |||
ZTF20aaqtncr | AT2021fez | 513 | 44.7d | 1320 | 40.3 | ||||
ZTF19aailpwl | AT2019brs | 173 | 45.1 | 2650 | 1000e | 40.4 | |||
ZTF20acvfraq | AT2020adpi | 120 | 44.6 | 3830 | 537 | 39.9 | |||
ZTF19aatubsj | AT2019fdr | 77 | 44.6 | 4100f | 370 | 40.1 | |||
ZTF20aanxcpf | AT2021loi | 59 | 44.0 | 2150 | 39.5 |
a) Revised from Wiseman et al. (2023) based on higher quality data. An additional broad component with FWHM km s-1 is also compatible with the data; b) the broad line is highly asymmetric and affected by imperfect telluric subtraction; c) there may be an even broader component with FWHM km s-1 but this may be caused by unmodelled Fe ii emission; d) Formally unconstrained due to only two available filters constrain blackbody model; e) consistent with the pre-flare measurement from an SDSS spectrum (Frederick et al., 2021); g) time of optical peak is uncertain; f) the significance of this component is dependent upon the continuum modelling
@@
ZTF20abrbeie | AT2021lwx | 113 | 113 | 54 | 54 | 45.34 | 3.54 | 0.04 | 0.06 | - |
ZTF19aamrjar | - | 25 | 25 | 44.55 | 3.28 | 0.09 | 0.09 | 44.9 | ||
ZTF20abodaps | AT2020afep | 220 | 220 | 44.69 | 3.20 | 0.15 | 0.15 | 44.6 | ||
ZTF18aczpgwm | AT2019kn | 339a | 592a | 77 | 77 | 44.42 | 3.30 | 0.07 | 0.07 | 44.4 |
@@
ZTF20aaqtncr | AT2021fez | 131 | 131 | 44.12 | 3.19 | 0.09 | 0.09 | 44.0 | ||
ZTF19aailpwl | AT2019brs | 331 | 389 | 116 | 116 | 44.41 | 3.20 | 0.11 | 0.11 | 44.3 |
ZTF20acvfraq | AT2020adpi | 160 | 160 | 44.18 | 3.12 | 0.12 | 0.12 | - | ||
ZTF19aatubsj | AT2019fdr | 303 | 303 | 82 | 82 | 44.00 | 3.08 | 0.12 | 0.12 | 44.6 |
ZTF20aanxcpf | AT2021loi | 174 | 365 | 70 | 70 | 43.59 | 3.14 | 0.06 | 0.06 | 43.1 |
@@ 3.2 Sample properties
In this section we present properties of the 10 selected ANTs measured from their light curves and spectra.
@@ 3.2.1 Black body properties
To estimate bolometric properties of the events, we model the observer-frame optical light curves as black bodies with evolving radii and temperatures using superbol (Nicholl, 2018)181818https://github.com/mnicholl/superbol. The resulting values are estimates that rely on several assumptions, namely that the emission is well described by a black body, in particular that it is spherically symmetric, and that thermal processes dominate over non-thermal emission. For three events, the presence of soft X-rays inconsistent with the UV-optical black body indicates either non-thermal emission or multiple black body emission regions. Likewise, the majority of the events show a delayed-onset mid-infrared (MIR) flare consistent with the dust echo mechanism (see Section 3.2.3). The events selected based on their durations and smoothness have peak black body luminosities in the range , and are thus among the most luminous transient astrophysical phenomena known, in particular more luminous than canonical TDEs (Yao et al., 2023a). We integrate the black body luminosities over the full observed light curve191919Note that some events are still ongoing, thus the energies are listed as lower limits to obtain the total radiated energy . The inferred energies are ten times greater than the most energetic supernovae (e.g. Nicholl et al., 2020) and three times the most energetic TDEs (Leloudas et al., 2016; Andreoni et al., 2022), and are firmly in the erg regime of the heterogeneously compiled ANTs in the literature (e.g. Leloudas et al., 2016; Graham et al., 2017; Kankare et al., 2017; Oates et al., 2024).
@@ 3.2.2 Spectroscopic properties
The ANT spectra are clearly distinct from those of TDEs, which are dominated by blue continua and low equivalent width broad emission lines - some show only H, while others show H+He, He alone, Bowen fluorescence lines and Fe, or are ‘featureless’ (Arcavi et al., 2014; Leloudas et al., 2019; van Velzen et al., 2021; Charalampopoulos et al., 2022). Our 10 ANTs are more AGN-like than TDE-like, consistent with the majority of ANTs in the literature. ANTs have narrower emission lines of higher equivalent width. Balmer lines of H are ubiquitous and have narrow and broad components, but the broad components are not as broad as those in TDEs: the mean H width from our 10 ANTs is km s-1 (Table 4) compared to km s-1 for TDEs (Charalampopoulos et al., 2022). Eight out of 10 ANTs have clear, narrow [O iii] which is a tracer of low-density ionized gas and present in both AGN and star-forming galaxy spectra (Vanden Berk et al., 2001; Osterbrock, 1989). It is lacking in the spectra of ZTF20abrbeie/AT2021lwx and ZTF20abodaps/AT2020afep. The ratio of this line to H can be used to estimate whether the ionizing source is an AGN or star formation (Baldwin et al., 1981; Kewley et al., 2001; Kauffmann et al., 2003) while its absence indicates that neither AGN emission or star formation are strong in the hosts of these transients.
@@ 3.2.3 Mid-infrared lags
All 10 events show changes in their MIR luminosity as measured in the WISE W1 (m) and W2 (m) bands. The light curve morphologies are varied, but all show a rise that is delayed from the optical flare. ZTF18aczpgwm, ZTF19aamrjar, ZTF20abodaps, ZTF20aaqtncr, and ZTF20aanxcpf show smooth, ongoing rises. On the other hand, ZTF20abrbeie, ZTF19aailpwl, ZTF20acvfraq and ZTF19aatubsj have MIR flares that peak within the timescale of the UV-optical light curve. MIR flares have been observed in TDEs (e.g. van Velzen et al., 2016; Jiang et al., 2021b), BFFs and ANTs (e.g. Jiang et al., 2021a; Hinkle, 2022; Wiseman et al., 2023) and are assumed to represent dust echoes from torus-like structures surrounding the black hole. The time lags between the UV-optical and MIR flares indicate the physical scale of the dust, the black body colour reveals its temperature, and the ratio of optical-to-MIR luminosity represents the dust covering fraction.
@@ 3.2.4 Dust properties
We estimate the dust temperature, radius and luminosity by fitting the host-subtracted and data with a black body SED. We infer a dust covering fraction as the ratio between the estimated dust luminosity and the UV-optical luminosity. Due to having only two WISE filters and no near IR, the blackbody temperatures are poorly constrained – formally the uncertainties are 100 per cent. Results are presented in Table 5. Our estimated dust properties are consistent with those measured for a sample of ANT MIR flares (Hinkle, 2022). In particular, we compare the three events common to both samples (ZTF19aatubsj, ZTF20cvfraq, ZTF20aanxcpf) for which our temperature and effective radius measurements agree.
@@
We perform a simple Kendall test for correlation for all pairs of parameters. In Table 6 we highlight all pairwise combinations for which the significance . The correlations between the and lags are expected, but we note that the ‘first’ lags, which are from the onset of the optical to the onset of the MIR, are not correlated with the peak lags. This is likely due to the 6-month NEOWISE cadence. We also recover expected correlations, albeit with fairly low significance, between the lags and the black body radius of the dust. The correlation between dust luminosity and temperature, and the UV-optical luminosity are further evidence that the MIR emission is indeed a dust echo of the UV-optical flare. The link between covering fraction and the lag, and dust radius, may indicate that larger dust structures have larger covering fractions, but may also be an artifact of fitting the dust black body with only two wavebands.
Interestingly the peak lag appears tentatively correlated with the width of narrow H but not with that of broad H. This may be evidence for the case where the narrow Balmer emission is caused by the reprocessing of kinetic energy of an outflow via the shock heating of a circumnuclear dust cloud, while the broad lines are from the outflow or accretion disk and thus do not correlate with the scale of the surrounding medium.
@@ 3.3 Summary of properties
In this section we have presented 10 examples of ANTs with yr light curves. Seven are photometrically selected, and three spectroscopically. They are luminous with erg s-1; the six ANTs from the photometric selection that have constraints on their black body SEDs have erg s-1. All ANTs have strong, high equivalent-width Balmer lines. These lines are narrow compared to TDEs. The mean FWHM of the broad component of H in the ANTs is km s-1 whereas for TDEs this is km s-1 (Charalampopoulos et al., 2022). All but one ANT has a strong MIR flare with a peak luminosity at least 10 per cent of the UV-optical black body luminosity, with temperature K, with radius pc and with a peak-to-peak lag from the optical d. All of these properties are consistent with a light echo from warm dust on the scales typically seen in AGN.

### v2

@@ 3.1.8 Spectroscopically selected ANTs
ZTF20acvfraq/AT2020adpi/ATLAS20bjzp/Gaia21aid, . First reported by ZTF via the AMPEL broker. The rise time is uncertain due to gaps in the observed light curve, constrained to d d in the rest frame. The initial decline is shallow, before a steep phase and then a second shallow phase.
We obtain a Keck/LRIS spectrum from WISeREP. The spectrum is similar to ZTf19aailpwl/AT2019brs: dominated by strong Balmer lines with broad (FWHM km s-1 and narrow components, along with He i and Mg ii with similar profiles. Similarly to ZTF19aailpwl/AT2019brs is the presence of He ii, N iii, and a particularly strong O iii line. These features indicate Bowen fluorescence. Fe ii features can also be seen in the spectrum.
A MIR flare is present with a lag of yr at peak. Swift observed the location of the transient over the period of a month around the optical peak.
Swift XRT observed this event eight times between MJD 59411 and 59456 ( d to d). In a total of ks, no X-rays are detected. Assuming a power-law with photon index 2 this corresponds to an upper limit of , which equates to a luminosity limit .
ZTF19aatubsj/AT2019fdr/ATLAS19lkd/PS19dar/Gaia19bsz, . First reported by ZTF via the AMPEL broker. Classified as a TDE in a NLSy1 (Frederick et al., 2021; Reusch et al., 2022), and although Pitik et al. (2022) classify it as a SLSN-IIn, we favour a SMBH-related interpretation.
The light curve is not smooth: the rise displays a ‘shoulder’ before peaking after 70 d, while after 70 d of steep decline there is an 80 d plateau before a shallower decline. After 500 days a 0.7 mag rebrightening occurred and the transient has remained close to that level ever since.
We retrieve the Keck/LRIS spectrum from WISeREP. The Balmer emission is the most prominent feature, and there is weak but clearly detected Heii.
There is a MIR flare with a lag at peak of d. The spectrum has narrow + broad Balmer emission lines as well as Fe ii emission. There is no clear detection of He ii or Bowen fluorescence lines. A MIR flare is present with a lag of yr at peak.
X-ray observations are presented in detail by Reusch
@@
ZTF20aanxcpf/AT2021loi/ATLAS21qje, . First reported by Gaia Science Alerts. This event was analysed by Makrygianni
et al. (2023) as a Bowen fluorescence flare. It is the second lowest luminosity of our sample ( erg s-1). It has a similar initial light curve shape to the majority of the analysis sample presented here, but displays a clear rebrightening after d in the rest frame.
The spectrum, from Keck/LRIS and obtained from WISeREP, is similar to ZTF19aailpwl/AT2019brs and ZTF20acvfraq/AT2020adpi in that it is dominated by Balmer lines with an unusual, flat-topped shape. There are strong emission lines from He ii, N iii and O iii, indicating Bowen fluorescence and thus a very hot ionizing source. Any Fe ii contribution is much weaker than in some of the Fe-strong events, but here the Balmer continuum is stronger.
A MIR flare is present with a lag of yr at peak.
Makrygianni
et al. (2023) report X-ray upper limits from Swift XRT of .
3.2 Unclear events
ZTF19aaozooc/AT2021hum/ATLAS20bjat, . The position of ZTF19aaozooc was targetted as part of the quasar survey (Richards et al., 2002) in the Sloan Digital Sky Survey (York et al., 2000). The spectrum, taken on MJD 52028 (19 years before the ZTF flare) and retrieved from SDSS data release 17 (Abdurro’uf et al., 2022), is of a type I quasar with broad Mg ii, C iii] and C iv. A further spectrum of ZTF19aaozooc was taken with Keck/LRIS around peak of the ZTF light curve (MJD 58788). The LRIS spectrum is qualitatively similar to the SDSS spectrum. We conclude that ZTF19aaozooc is a type I quasar with slow, high-amplitude variability.
ZTF20aaqtncr/AT2021fez/ATLAS20pzv/Gaia21bgs, . First reported by Gaia Science Alerts. The light curve has a smooth rise and decline of similar duration, although the ZTF data release photometry (without difference imaging) shows variability before the transient detection. A spectrum was obtained with LCO Floyds FTN. The spectrum shows strong narrow Balmer emission lines, strong [N ii], and very strong [O iii]. The [O iii]/H ratio alone is enough to place the ionizing source as an AGN, which is confirmed by placing the line fluxes in the [O iii]/H vs [N ii/H] Baldwin-Phillips-Terlevich (BPT) diagram (Baldwin
@@ 3.2 Unclear events
ZTF20abrbeie | AT2021lwx | 198 | 2560a | 250 | - | ||||
ZTF19aamrjar | - | 377 | 6200b | 732 | 41.5 | ||||
ZTF20abodaps | AT2020afep | 17 | 1540 | 40.7 | |||||
ZTF18aczpgwm | AT2019kn | 125 | 2831c | 245 | 40.6 | ||||
ZTF21abxowzx | AT2021yzu | 71 | 1690 | 194 | 40.2 | ||||
ZTF19aailpwl | AT2019brs | 173 | 2650 | 1000e | 40.4 | ||||
ZTF20abgxlut | AT2020oio | 96 | 3800 | 970 | 39.4 | ||||
ZTF22aadesap | AT2022pfx | 103 | 2100 | 39.4 | |||||
ZTF20acvfraq | AT2020adpi | 120 | 3830 | 537 | 39.9 | ||||
ZTF19aatubsj | AT2019fdr | 77 | 4100f | 370 | 40.1 | ||||
ZTF20aanxcpf | AT2021loi | 59 | 2150 | 39.5 |
a) Revised from Wiseman et al. (2023) based on higher quality data. An additional broad component with FWHM km s-1 is also compatible with the data; b) the broad line is highly asymmetric and affected by imperfect telluric subtraction; c) there may be an even broader component with FWHM km s-1 but this may be caused by unmodelled Fe ii emission; d) Formally unconstrained due to only two available filters constrain blackbody model; e) consistent with the pre-flare measurement from an SDSS spectrum (Frederick et al., 2021); f) the significance of this component is dependent upon the continuum modelling; g) using only as opposed to Koljonen et al. (2024) who add Swift UV data.
@@
ZTF20abrbeie | AT2021lwx | 119 | 119 | 54 | 54 | 44.63 | 3.54 | 0.04 | 0.06 | - |
ZTF19aamrjar | - | 472 | 472 | 25 | 25 | 44.55 | 3.28 | 0.09 | 0.09 | 44.9 |
ZTF20abodaps | AT2020afep | 220 | 220 | 44.69 | 3.20 | 0.15 | 0.16 | 44.6 | ||
ZTF18aczpgwm | AT2019kn | 77 | 77 | 44.42 | 3.30 | 0.07 | 0.17 | 44.4 | ||
@@
ZTF19aailpwl | AT2019brs | 405 | 405 | 116 | 116 | 44.41 | 3.20 | 0.11 | 0.20 | 44.3 |
ZTF20abgxlut | AT2020oio | 139 | 139 | 43 | 43 | 44.13 | 3.16 | 0.09 | 0.86 | - |
ZTF22aadesap | AT2022pfx | 137 | 286 | 260 | 69 | 43.02 | 3.02 | 0.05 | 0.18 | 43.5 |
ZTF20acvfraq | AT2020adpi | 160 | 160 | 44.18 | 3.12 | 0.12 | 0.38 | - | ||
ZTF19aatubsj | AT2019fdr | 303 | 303 | 82 | 82 | 44.00 | 3.08 | 0.12 | 0.25 | 44.6 |
ZTF20aanxcpf | AT2021loi | 173 | 364 | 70 | 70 | 43.59 | 3.14 | 0.06 | 0.39 | 43.1 |
@@ 3.3 Sample properties
In this section we present properties of the 11 selected ANTs measured from their light curves and spectra.
@@ 3.3.1 Black body properties
To estimate bolometric properties of the events, we model the observer-frame optical light curves as black bodies with evolving radii and temperatures using superbol (Nicholl, 2018)202020https://github.com/mnicholl/superbol. The resulting values are estimates that rely on several assumptions, namely that the emission is well described by a black body, in particular that it is spherically symmetric, and that thermal processes dominate over non-thermal emission. For three events, the presence of soft X-rays inconsistent with the UV-optical black body indicates either non-thermal emission or multiple black body emission regions. Likewise, the majority of the events show a delayed-onset mid-infrared (MIR) flare consistent with the dust echo mechanism (see Section 3.3.3). Black body measurements are presented in Table 4. The ANTs have peak black body luminosities in the range , and are thus among the most luminous transient astrophysical phenomena known, in particular more luminous than canonical TDEs (Yao et al., 2023a). We integrate the black body luminosities over the full observed light curve212121Note that some events are still ongoing, thus the energies are listed as lower limits to obtain the total radiated energy . The inferred energies are ten times greater than the most energetic supernovae (e.g. Nicholl et al., 2020) and three times the most energetic TDEs (Leloudas et al., 2016; Andreoni et al., 2022), and are firmly in the erg regime of the heterogeneously compiled ANTs in the literature (e.g. Leloudas et al., 2016; Graham et al., 2017; Kankare et al., 2017; Oates et al., 2024).
@@ 3.3.2 Spectroscopic properties
The ANT spectra are clearly distinct from those of TDEs, which are dominated by blue continua and low equivalent width broad emission lines - some show only H, while others show H+He, He alone, Bowen fluorescence lines and Fe, or are ‘featureless’ (Arcavi et al., 2014; Leloudas et al., 2019; van Velzen et al., 2021; Charalampopoulos et al., 2022). Our 11 ANTs are more AGN-like than TDE-like, consistent with the majority of ANTs in the literature. ANTs have narrower emission lines of higher equivalent width. Balmer lines of H are ubiquitous and have narrow and broad components, but the broad components are not as broad as those in TDEs: the mean H width from our 11 ANTs is km s-1 (Table 4) compared to km s-1 for TDEs (Charalampopoulos et al., 2022). Eight out of 11 ANTs have clear, narrow [O iii] which is a tracer of low-density ionized gas and present in both AGN and star-forming galaxy spectra (Vanden Berk et al., 2001; Osterbrock, 1989). It is lacking in the spectra of ZTF20abrbeie/AT2021lwx and ZTF20abodaps/AT2020afep. The ratio of this line to H can be used to estimate whether the ionizing source is an AGN or star formation (Baldwin et al., 1981; Kewley et al., 2001; Kauffmann et al., 2003) while its absence indicates that neither AGN emission or star formation are strong in the hosts of these transients.
@@ 3.3.3 Mid-infrared lags
All 11events show changes in their MIR luminosity as measured in the WISE W1 (m) and W2 (m) bands. The light curve morphologies are varied, but all show a rise that is delayed from the optical flare. ZTF18aczpgwm, ZTF19aamrjar, ZTF20abodaps, ZTF20aaqtncr, and ZTF20aanxcpf show smooth, ongoing rises. On the other hand, ZTF20abrbeie, ZTF19aailpwl, ZTF20acvfraq and ZTF19aatubsj have MIR flares that peak within the timescale of the UV-optical light curve. MIR flares have been observed in TDEs (e.g. van Velzen et al., 2016; Jiang et al., 2021b), BFFs and ANTs (e.g. Jiang et al., 2021a; Hinkle, 2024; Wiseman et al., 2023) and are assumed to represent dust echoes from torus-like structures surrounding the black hole. The time lags between the UV-optical and MIR flares indicate the physical scale of the dust, the black body colour reveals its temperature, and the ratio of optical-to-MIR luminosity represents the dust covering fraction.
@@ 3.3.4 Dust properties
We estimate the dust temperature, radius and luminosity by fitting the host-subtracted and data with a black body SED. We infer a dust covering fraction as the ratio between the estimated dust luminosity and the UV-optical luminosity. Due to having only two WISE filters and no near IR, the blackbody temperatures are poorly constrained – formally the uncertainties are 100 per cent. Results are presented in Table 5. Our estimated dust properties are consistent with those measured for a sample of ANT MIR flares (Hinkle, 2024). In particular, we compare the three events common to both samples (ZTF19aatubsj, ZTF20cvfraq, ZTF20aanxcpf) for which our temperature and effective radius measurements agree. The covering fractions are of similar magnitude to those reported by Hinkle (2024). We note that dust emission is not a perfect black body: the main effects of allowing for an absorption coefficient would be a shift to lower inferred dust temperatures and larger distances from the source to the dust (Jiang et al., 2021b), although neither of these affect our conclusions.
@@
We perform a simple Kendall test for correlation for all pairs of parameters. To determine which correlations are significant, we check for those with a -value of 0.05 adapted for the look elsewhere effect for 50 parameter pairs, resulting in a -value threshold of 0.001. Three pairs pass this threshold: the ‘first’ lags and peak lags, although we note that these two lag measurements are not correlated with each other. This is likely due to the 6-month NEOWISE cadence. We also recover a correlation between dust luminosity the UV-optical luminosity ( ), further evidence that the MIR emission is indeed a dust echo of the UV-optical flare.
We also note that the and peak lags appear tentatively correlated with the width of narrow H ( , with the high -value due to some events not having a clear narrow feature) but not with that of broad H. This may be evidence for the case where the narrow Balmer emission is caused by the reprocessing of kinetic energy of an outflow via the shock heating of a circumnuclear dust cloud, while the broad lines are from the outflow or accretion disk and thus do not correlate with the scale of the surrounding medium.
@@ 3.4 Summary of properties
In this section we have presented 11 examples of ANTs with yr light curves. Seven are photometrically selected, and three spectroscopically. They are luminous with erg s-1; the six ANTs from the photometric selection that have constraints on their black body SEDs have erg s-1. All ANTs have strong, high equivalent-width Balmer lines. These lines are narrow compared to TDEs. The mean FWHM of the broad component of H in the ANTs is km s-1 whereas for TDEs this is km s-1 (Charalampopoulos et al., 2022). All but one ANT has a strong MIR flare with a peak luminosity at least 10 per cent of the UV-optical black body luminosity, with temperature K, with radius pc and with a peak-to-peak lag from the optical d. All of these properties are consistent with a light echo from warm dust on the scales typically seen in AGN.